OpenStudy (anonymous):

if vector a=(-4,5) and vector b=(2-8) what is ||a-b|| and ||a||-||b|| PLEASE HELP!

jimthompson5910 (jim_thompson5910):

are you able to determine what `a-b` is equal to?

OpenStudy (anonymous):

yes its (-6, 13)

OpenStudy (anonymous):

the answer for ||a-b|| is (-8,10) but I DONT KNOW HOW TO GET THERE

jimthompson5910 (jim_thompson5910):

||a-b|| should be a single number

jimthompson5910 (jim_thompson5910):

||a-b|| represents the length or magnitude of vector `a-b`

jimthompson5910 (jim_thompson5910):

if a=(-4,5) b=(2,-8) then a-b = (-6, 13) I agree with what you got for `a-b`

jimthompson5910 (jim_thompson5910):

now you use this formula if z = (m,n) then \[\Large ||z|| = \sqrt{m^2 + n^2}\]

OpenStudy (anonymous):

Yes I know thats why I got so confused, cause I know that a Magnitude is just a number, but the answer is a coordinate so I don't know what to do or how o get to those coordinates

jimthompson5910 (jim_thompson5910):

I'm not sure what they're really asking. Perhaps it's a typo or mixup?

OpenStudy (superdavesuper):

@nicolepenafielm are u sure "the answer is a coordinate"? The Q asks "what is ||a-b|| and ||a||-||b||" so it could be TWO separate nos even thou they look to be coordinates...

OpenStudy (anonymous):

Number 12 is the exercise! please help

OpenStudy (anonymous):

sorry is number 11

OpenStudy (superdavesuper):

@nicolepenafielm ur screen shot shows the Q asking to sketch the vectors listed - its completely different from what u posted about | |....

jimthompson5910 (jim_thompson5910):

if you're talking about 11c, then a = (-4,5) 2a = 2*(-4,5) 2a = (2*(-4), 2*5) 2a = (-8, 10)

jimthompson5910 (jim_thompson5910):

I think you mixed up 11c and 12c somehow

OpenStudy (anonymous):

you are completely right sorry

OpenStudy (anonymous):

Linear algebra?